2.1 Suppose the market model is described by the following demand and supply functions: Qd= 200-2P+ 4i and Qs= -200 + 3P where i =50 Answer the following questions: a) Calculate the equilibrium price (P) and quantity (Q) and indicate your answers on the graph. [03] b) If the market demand decreases by 50% what will be the new clearing price and quantity for this market? [03] c) Indicate the effect of the changes in (b) on a graph and explain what will happen to market price and quantity. [03]


1
Expert's answer
2022-02-15T04:20:52-0500

"Q_d= 200-2P+4i"

i= 50

"Q_d= 400-2P"

"Q_s= -200+3P"


a) At Equilibrium, "Q_s=Q_d"

"400-2P=-200+3P"

"600= 5P"

P= 120

Q*= -200+ 3(120)

= -200+ 360

= 160


b) A decrease in Demand by 50%

New Demand Function= 0.5( 400- 2P)

"Q_d= 200-P"

New Equilibrium Price and Quantity

200-P= -200+3P

400=4P

P= 100

Q*= 200- 100

= 100

c) Decrease in demand causes a decline in the and also the quantity demanded.

This can be shown in the diagram below;